OpenStudy (anonymous):

Find the integral: 1/(y-y^2)dy

OpenStudy (anonymous):

I 'm pretty sure you separate into 1/y and 1/(1-y), then using integration by parts, but no matter which variable I choose for u and v, the problem only gets uglier. \[\int\limits1/(y-y^2)dy\]

OpenStudy (anonymous):

I don't think you need to use integration by parts to solve this. Using partial fractions, you can get 2 integrals: \[\int\limits_{}^{} 1/y\ dy + \int\limits_{}^{} 1/(1-y)\ dy\] Partial Fractions: \[1/y(1-y) = A/y + B/(1-y)\] \[1=A(1-y) + By\] \[1=(B-A)y+A\] A & B both equal 1. So now integrate those 2 integrals and see what you get. I got: \[\ln \left| y \right| - \ln \left| 1-y \right| +C\] There is a minus sign before the 2nd ln because when you take the integral of 1/(1-y) and substitute u=(1-y), du=(-1), so you get -ln.

OpenStudy (anonymous):

I'm thinking you're correct, but I've got to give it a second for my brain to make sense of it... :-p

OpenStudy (anonymous):

I had to do it a few times and reference a couple different sources too; it's the coefficients part of the partial fractions that was tripping me up :P Try wolframalpha.com if you don't already make use of it :)

OpenStudy (anonymous):

That's a huge help, thanks! I've been fighting this problem for about an hour.

OpenStudy (anonymous):

Too bad that was only 1 tiny part of the problem I'm dealing with! lol! !&(%^(& differential equations.

OpenStudy (anonymous):

I know... I always seem to manage to drop negative signs & powers mysteriously "disappear" because my brain feels like mush. I'm about to review integration techniques for my exam next week: trig substitution and partial fractions :P I'll tackle improper integrals and area/volume problems tomorrow... hopefully, haha. Good luck!

OpenStudy (anonymous):

You too, thanks for the help.